# Polarization Leakage and the IXPE PSF

### v1

@@ 2.1 Polarization Leakage
We start by Taylor expanding the integrand of Eq. 1 in $\bm{\delta}$ to fourth order, keeping only the even powers of $\bm{\delta}$. While $P(\bm{\delta}|\phi)$ is asymmetric due to conversion point uncertainty associated with the Bragg peak, especially for Mom reconstruction, the polarization is intrinsically invariant under ${}^{\circ}$ rotations. Odd orders of $\bm{\delta}$ flip sign under these rotations, so they integrate out and we are left with (in index notation)
@@
For an unpolarized source ($\Pi=0$), Eq. 2.1 predicts
@@
where $\sigma_{\pm}^{2}=\sigma_{\parallel}^{2}\pm\sigma_{\perp}^{2}$ and $k_{\pm}^{4}=k_{\parallel}^{4}\pm k_{\perp}^{4}$. These are the contribution of each photoelectron to $I$, $Q$, and $U$ maps. The total maps are exactly the average of these functions over all photoelectrons, which is equal to these functions applied with the parameters of the average photoelectron due to the following argument. Eqs. 2.1 are photoelectron-independent except for the energy dependence of $\sigma_{\parallel}$, $\sigma_{\perp}$, and $k_{\parallel}$. Since Eqs. 2.1 are linear in the energy dependent terms, the $I$, $Q$, and $U$ maps obey superposition. Taking $\sigma_{\parallel}^{2}$ to be the spectrum-weighted averages of $\sigma^{2}_{\parallel}(E)$ and so on for the other parameters therefore produces the ensemble map. We neglect the energy dependence of the mirror PSF, which becomes important only above 4–5 keV and would violate this argument.
Qualitatively, Eqs. 2.1 demonstrate the effect of polarization leakage. There is an isotropic blur on $I$, equivalent, to first order, to a Gaussian blur with width equal to the quadrature mean of $\sigma_{\parallel}$ and $\sigma_{\perp}$. For $\sigma_{\parallel}\neq\sigma_{\perp}$, there is also a leakage pattern in $Q$ and $U$ that depends on the derivatives of the PSF. These effects occur even if the source is unpolarized.
@@
have also been used. To produce maps, the photoelectron weight $\mu$ must be averaged similarly to $\sigma_{\parallel}$ above. These terms effectively bias the blurring of $I$ according to the source polarization (Eq. 7a) and add a PSF-distributed source $Q$ and $U$ map to the leakage pattern with additional corrections (Eqs. 7b, 7c). For connection with Bucciantini et al. (2023), Eqs. 2.1 are also expressed in terms of Mueller matrices in appendix B.
Notice that the source polarization effects (Eqs. 2.1) are reduced by $\mu<1$ while source polarization-independent leakage (Eqs. 2.1) is not. This highlights the importance of leakage correction, especially when source polarizations are low.
@@ 2.3 Sky-calibrated PSFs
Our fit method uses gradient descent to minimize the $\chi^{2}$ value comparing the blurred PSF predicted by Eqs. 2.1 in each detector to observations. We reuse the regulatory term defined in Eq. 11 to prevent the amplification of noise. The blurred PSFs—and their $\chi^{2}$ statistic when matched to data—are analytical functions of the PSF images and can be differentiated by hand as in the previous section.
@@ 3.3 Polarization mapping with leakage/PSF models
[…]
@@ Appendix B Mueller Matrices
    $\left(\begin{matrix}I(\bm{r})\\ Q(\bm{r})\\ U(\bm{r})\end{matrix}\right)=M(\bm{r}-\bm{r}^{\prime})\left(\begin{matrix}I_{\mathrm{src}}(\bm{r}^{\prime})\\ Q_{\mathrm{src}}(\bm{r}^{\prime})\\ U_{\mathrm{src}}(\bm{r}^{\prime})\end{matrix}\right)$ (B1)
    for Mueller matrix $M(\bm{r}-\bm{r}^{\prime})$. Eqs. 2.1 for an unpolarized point source dictates that
@@
    using functions defined in 2.1 and 2.1. Note that $M(\bm{x})$ can be computed directly from the PSF without any prior knowledge of the source aside from the parameters $\sigma_{\parallel}$, $\sigma_{\perp}$, $k_{\parallel}$, and $\mu$, which depend on the source spectrum. Polarization maps can be computed for an extended source by convolving $M$ with the source polarization maps—i.e., by integrating the left hand side of B1 with respect to $\bm{r}^{\prime}$.

### v2

@@ 2.1 Polarization Leakage
We start by Taylor expanding the integrand of Eq. 1 in $\bm{\delta}$ to fourth order, keeping only the even powers of $\bm{\delta}$. While $P(\bm{\delta}|\phi)$ is asymmetric due to conversion point uncertainty associated with the Bragg peak, especially for Mom reconstruction, the polarization is intrinsically invariant under 180 rotations. Odd orders of $\bm{\delta}$ flip sign under these rotations, so they integrate out and we are left with (in index notation)
@@
For an unpolarized source ($\Pi=0$), Eq. 2 predicts
@@
where $\sigma_{\pm}^{2}=\sigma_{\parallel}^{2}\pm\sigma_{\perp}^{2}$ and $k_{\pm}^{4}=k_{\parallel}^{4}\pm k_{\perp}^{4}$. These are the contribution of each photoelectron to $I$, $Q$, and $U$ maps. The total maps are exactly the average of these functions over all photoelectrons, which is equal to these functions applied with the parameters of the average photoelectron due to the following argument. Eqs. 6 are photoelectron-independent except for the energy dependence of $\sigma_{\parallel}$, $\sigma_{\perp}$, and $k_{\parallel}$. Since Eqs. 6 are linear in the energy dependent terms, the $I$, $Q$, and $U$ maps obey superposition. Taking $\sigma_{\parallel}^{2}$ to be the spectrum-weighted averages of $\sigma^{2}_{\parallel}(E)$ and so on for the other parameters therefore produces the ensemble map. We neglect the energy dependence of the mirror PSF, which becomes important only above 4–5 keV and would violate this argument.
Qualitatively, Eqs. 6 demonstrate the effect of polarization leakage. There is an isotropic blur on $I$, equivalent, to first order, to a Gaussian blur with width equal to the quadrature mean of $\sigma_{\parallel}$ and $\sigma_{\perp}$. For $\sigma_{\parallel}\neq\sigma_{\perp}$, there is also a leakage pattern in $Q$ and $U$ that depends on the derivatives of the PSF. These effects occur even if the source is unpolarized.
@@
have also been used. To produce maps, the photoelectron weight $\mu$ must be averaged similarly to $\sigma_{\parallel}$ above. These terms effectively bias the blurring of $I$ according to the source polarization (Eq. 7a) and add a PSF-distributed source $Q$ and $U$ map to the leakage pattern with additional corrections (Eqs. 7b, 7c). For connection with Bucciantini et al. (2023), Eqs. 7 are also expressed in terms of Mueller matrices in appendix B.
Notice that the source polarization effects (Eqs. 7) are reduced by $\mu<1$ while source polarization-independent leakage (Eqs. 6) is not. This highlights the importance of leakage correction, especially when source polarizations are low.
@@ 2.3 Sky-calibrated PSFs
Our fit method uses gradient descent to minimize the $\chi^{2}$ value comparing the blurred PSF predicted by Eqs. 6 in each detector to observations. We reuse the regulatory term defined in Eq. 11 to prevent the amplification of noise. The blurred PSFs—and their $\chi^{2}$ statistic when matched to data—are analytical functions of the PSF images and can be differentiated by hand as in the previous section.
@@ 3.3 Polarization mapping with leakage/PSF models
The method successfully recovers small polarization structures within the nebula. The broad strokes of the NASA text, the red ribbon, the base of the Stanford tree, and the S are all recovered despite not appearing in the raw observations. All of these structures have widths below that of any of the PSFs. However, leakage from the central point source obscures nearby nebula polarization. The maps extracted using the symmetric PSFs (3rd column) show the risks of using an inaccurate model. All the major structures of the nebula are recovered, but the leakage of the central point source is incompletely subtracted leaving incorrect nearby polarization predictions. Far from the point source, the image is blotchier and sharp edges in the true polarization map are incorrectly localized.
@@ Acknowledgments
Appendix A Auxiliary Leakage Matrices
We defined eight matrices to aid in computing the derivatives of the PSF $P_{\mathrm{mir}}(\bm{r})$ in section 2.1. They are
$\begin{split}\mathsf{Z}_{\sigma}=\frac{1}{(4\Delta)^{2}}\left(\begin{array}[]{ccccccccc}\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&1&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&1&\cdot&-4&\cdot&1&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&1&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \end{array}\right)\qquad&\mathsf{Z}_{k}=\frac{1}{4(4\Delta)^{4}}\left(\begin{array}[]{ccccccccc}\cdot&\cdot&\cdot&\cdot&1&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&2&\cdot&-8&\cdot&2&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ 1&\cdot&-8&\cdot&20&\cdot&-8&\cdot&1\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&2&\cdot&-8&\cdot&2&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&1&\cdot&\cdot&\cdot&\cdot\\ \end{array}\right)\\ \mathsf{Q}_{\sigma}=\frac{1}{(4\Delta)^{2}}\left(\begin{array}[]{ccccccccc}\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&1&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&-1&\cdot&\cdot&\cdot&-1&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&1&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \end{array}\right)\qquad&\mathsf{Q}_{k}=\frac{1}{3(4\Delta)^{4}}\left(\begin{array}[]{ccccccccc}\cdot&\cdot&\cdot&\cdot&1&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&-4&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ -1&\cdot&4&\cdot&\cdot&\cdot&4&\cdot&-1\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&-4&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&1&\cdot&\cdot&\cdot&\cdot\\ \end{array}\right)\\ \mathsf{U}_{\sigma}=\frac{1}{(4\Delta)^{2}}\left(\begin{array}[]{ccccccccc}\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&-2&\cdot&2&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&2&\cdot&-2&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \end{array}\right)\qquad&\mathsf{U}_{k}=\frac{1}{3(4\Delta)^{4}}\left(\begin{array}[]{ccccccccc}\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&-2&\cdot&2&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&-2&\cdot&12&\cdot&-12&\cdot&2&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&2&\cdot&-12&\cdot&12&\cdot&-2&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&2&\cdot&-2&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \end{array}\right)\\ \mathsf{X}_{k}=\frac{1}{3(4\Delta)^{4}}\left(\begin{array}[]{ccccccccc}\cdot&\cdot&\cdot&\cdot&-1&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&6&\cdot&-8&\cdot&6&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ -1&\cdot&-8&\cdot&12&\cdot&-8&\cdot&-1\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&6&\cdot&-8&\cdot&6&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&-1&\cdot&\cdot&\cdot&\cdot\\ \end{array}\right)\qquad&\mathsf{Y}_{k}=\frac{1}{3(4\Delta)^{4}}\left(\begin{array}[]{ccccccccc}\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&4&\cdot&-4&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&-4&\cdot&\cdot&\cdot&\cdot&\cdot&4&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&4&\cdot&\cdot&\cdot&\cdot&\cdot&-4&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&-4&\cdot&4&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot&\cdot\\ \end{array}\right)\end{split}$ (A1)
where $\Delta$ is the width of pixels in the PSF image and dots indicate zeros.
@@ Appendix B Mueller Matrices
$\left(\begin{matrix}I(\bm{r})\\ Q(\bm{r})\\ U(\bm{r})\end{matrix}\right)=\int d^{2}\bm{r}^{\prime}\,M(\bm{r}-\bm{r}^{\prime})\left(\begin{matrix}I_{\mathrm{src}}(\bm{r}^{\prime})\\ Q_{\mathrm{src}}(\bm{r}^{\prime})\\ U_{\mathrm{src}}(\bm{r}^{\prime})\end{matrix}\right)$ (B1)
for Mueller matrix $M(\bm{r}-\bm{r}^{\prime})$. Eqs. 7 for an unpolarized point source dictates that
@@
using functions defined in 6 and 8. Note that $M(\bm{x})$ can be computed directly from the PSF without any prior knowledge of the source aside from the parameters $\sigma_{\parallel}$, $\sigma_{\perp}$, $k_{\parallel}$, and $\mu$, which depend on the source spectrum. Polarization maps can be computed for an extended source by convolving $M$ with the source polarization maps—i.e., by integrating the left hand side of B1 with respect to $\bm{r}^{\prime}$.